OpenStudy (anonymous):

There is a formula used in chemistry and physics that states that PV = nrt. Solve this formula for n.

OpenStudy (whpalmer4):

\[PV = nrt\]If you want to solve that for \(n\), remember that you can multiply or divide the whole equation by any number or expression as long as you do it to both sides. For example, if we divide both sides by \(r\) we get \[\frac{PV}r = \frac{nrt}r\]or\[\frac{PV}r = nt\]after canceling \(r\) from both numerator and denominator on the right. Continue the process to solve for \(n\).

OpenStudy (aaronq):

T and R should be capitalized ;)